If Emma is 5 years older than 6 times Josh's age. And Emma is 29 years old.
Answer by mananth(16946)

Let Josh be = x years
6 times x = 6x
5 more than 6x=6x+5
Emma= 6x+5
6x+5 =29
..
-5 -5
6x=29-5
6x=24
/6
x= 4 years Josh's age
..
m.ananth@hotmail.ca